OpenStudy (anonymous):

WILL MEDAL!!! Which of the following equations is equivalent to log(y)=3.994 ? A.y=10+3.994^3.994 B.y=3.994^10 C.y=3.994^3.994 D.y=10^3.994

OpenStudy (trojanpoem):

loga = b 10^b = a in your case 10^3.994 = y

OpenStudy (anonymous):

i know nothing about prob and stats

OpenStudy (anonymous):

10^3.994=y